a sphere of mass m1=2kg collides with a sphere of mass m2=3kg which is at rest.mass m1 will move at right angies to the line,joining centres at the time of collision,if the coefficient of restitution is


Grade:10

1 Answers

Aman singh
5 years ago
  1. Apply conservation of linear momentum about line of impact taking theta angle from velocity and LOI so,  we get m1ucosΦ=m2v2
  2. And by newtons law of restitution we get, v2/ucosΦ=e and v2/ucosΦ is equal to m1/m2 in above equation so e=m1/m2
  3. e=2/3